OpenStudy (babybritt002):

Twice Jack's age plus 5 is 21. Write and solve an equation. How old is Jack?

OpenStudy (calculusxy):

Well set up an equation again: 2x + 5 = 21 Solve for x now.

OpenStudy (anonymous):

jack can be 8 because once you multiply that by 2 you get 16 and 16 plus 5 is 21 (:

OpenStudy (anonymous):

in that's exactly what you have in the equation lol , just saw that. @calculusxy

OpenStudy (calculusxy):

Yes. Look at the equation: 2x + 5 = 21 21 - 5 = 16 2x = 16 Divide both sides by 2 16/2=8 x = 8 Proof: 2(8)=16 + 5 = 21